OpenStudy (anonymous):

Will Fan and Medal! Thank you♥ 1) Identify the function rule shown in the table. Function Table n - 3, 4, 5, 6 y - 2, 1, 0, -1 a. y = 2 + n b. y = 5n c. y = 5 - n d. not enough information

OpenStudy (anonymous):

@flower2468

OpenStudy (flower2468):

okay so for a 3+2=5 and there is no 5 in y so not a for b 3*5=15 and there is no 15 in y so not b for c 5-3=2 and there is a 2 in y, but to make sure we move to the next number in line 5-4=1 and 1 is in they line so C is correct

OpenStudy (anonymous):

@flower2468 Thank you so much♥

OpenStudy (flower2468):

you are welcome:)